OpenStudy (anonymous):

Solve by completing the square: x^2-8x+13=0 My answer was x=4 +/- sqrt. 29 Please check my work, correct me if i'm wrong!

OpenStudy (anonymous):

(x-4)^2 + 13 -16 = 0 (x-4)^2 = 3 x-4 = +- sqrt3 x = 4 + sqrt3 or 4 - sqrt3

OpenStudy (anonymous):

i think you got your signs mixed up 13+16 = 29 its -13 + 16 on RHs making 3

OpenStudy (anonymous):

oh, okay thanks! :D

OpenStudy (anonymous):

no probs